OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side

OpenStudy (anonymous):

5n = n+16 5n/n = 16 4n = 16 n = 16/4 n = 4 Side length = 5n = 5*4 = 20